The square of the sum of a number and 2 is thrice the difference of the number and 2. Find the numbers.
Answer by ankor@dixie-net.com(22740)

Write an equation for what it says
:
"The square of the sum of a number and 2 is thrice the difference of the number and 2."
(n+2)^2 = 3(n-2)
FOIL (n+2)(n+2). distribute
n^2 + 2n + 2n + 4 = 3n - 6
Combine like terms on the left
n^2 + 4n - 3n + 4 + 6 = 0
n^2 + n + 10 = 0
This equation has no real solution, so the answer is no solution
:
Tried this with a right side written like this
(n+2)^2 = 3(2-n)
FOIL (n+2)(n+2). distribute
n^2 + 2n + 2n + 4 = 6 - 3n
Combine like terms on the left
n^2 + 4n + 3n + 4 - 6 = 0
n^2 + 7n - 2 = 0
here we got solutions
n = -7.275
n = .275
These solutions did work in the original equation